# Hausdorffness of certain nilpotent cohomology spaces

### v1

@@ Abstract.
Mathematics Subject Index Mathematics Subject Classification:
@@ 1. Introduction
Let $G$ be a Lie group. In our context, a representation of $G$ is a quasi-complete, locally convex, Hausdorff complex topological space $V$, together with a continuous linear action
@@
Let $\mathfrak{g}$ be a finite-dimensional complex Lie algebra. A continuous $\mathfrak{g}$-module is a quasi-complete, locally convex, Hausdorff complex topological space $V$, together with a continuous Lie algebra action
@@
Acknowledgment: The authors would like to thank Wei Xiao for suggesting the references [5, 6, 8] on the cubic Dirac operator. F. J. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the grant SFB-TRR 358/1 2023-491392403. B. Sun was supported in part by National Key R & D Program of China (Nos. 2022YFA1005300 and 2020YFA0712600) and New Cornerstone Investigator Program.
@@ 4. Proof of the main theorem
We keep the notations and assumptions from previous sections. In particular, $G$ is a (possibly disconnected) compact Lie group with complexified Lie algebra $\mathfrak{g}$, $(\pi,V)$ is a smooth representation of $G$ and $\mathfrak{q}$ is parabolic subalgebra of $\mathfrak{g}$ with nilpotent radical $\mathfrak{u}$. Use the notations in Section 3, $S=\wedge^{\bullet}\mathfrak{u}$, $\mathfrak{h}$ is a common Cartan subalgebra of $\mathfrak{l}$ and $\mathfrak{g}$, $D$ is the cubic Dirac operator with respect to $(\mathfrak{g},\mathfrak{l})$ and $\partial\in{\mathrm{End}}(V\otimes S)$ (resp. $\operatorname{d}\in{\mathrm{End}}(V\otimes S)$) are the boundary (resp. coboundary) map of Lie algebra homology (resp. cohomology). From now on, we view $D$ as an operator on $V\otimes S$.
@@
Let $\mathfrak{b}_{L}$ be the Borel subalgebra of $\mathfrak{l}$ associated with $\Delta_{\mathfrak{l}}^{+}$ and $T_{L}:=N_{L}(\mathfrak{b}_{L})$ be the Cartan subgroup of $L$. Recall that $T_{L}$ acts on $\mathfrak{h}^{*}$ by the coadjoint action, which preserves $\Delta_{\mathfrak{l}}^{+}$, and inner form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is invariant under the coadjoint action. By the Cartan-Weyl theory (see [7, Chapter 4, Section 2]), there is a finite set $\{t_{1},\cdots,t_{k}\}$ of elements in the Cartan subgroup $T_{L}$ of $L$, such that as $\mathfrak{l}$-modules,

### v2

@@ Abstract.
2020 Mathematics Subject Classification:
@@ 1. Introduction
Let $G$ be a Lie group. In our context, a representation of $G$ is a quasi-complete, locally convex, Hausdorff complex topological vector space $V$, together with a continuous linear action
@@
Let $\mathfrak{g}$ be a finite-dimensional complex Lie algebra. A continuous $\mathfrak{g}$-module is a quasi-complete, locally convex, Hausdorff complex topological vector space $V$, together with a continuous Lie algebra action
@@
Our results are related to Vogan’s conjecture through Hochschild-Serre spectral sequence, which is one motivation of our work. Recall that $V$ is a smooth representation of a real reductive group $G$. Note that $\mathfrak{u}\cap\mathfrak{k}$ is a nilpotent subalgebra of $\mathfrak{k}$, where $\mathfrak{k}=\mathfrak{g}^{\theta}$ is the complexified Lie algebra of $K$.
Ignoring the topology, there exists a convergent spectral sequence $\{E_{r}^{p,q}\}_{r\geq 0}\Rightarrow{\mathrm{H}}^{p+q}(\mathfrak{u},V)$ in the category of $\mathfrak{l}\cap\mathfrak{k}$-modules, which is called the Hochschild-Serre spectral sequence. The $E_{1}$-terms are given by $E_{1}^{p,q}={\mathrm{H}}^{n(p,q)}(\mathfrak{u}\cap\mathfrak{k},V)\otimes X_{p}$. Here $X_{p}$ is a finite-dimensional vector space and $n(p,q)$ is a certain integer depending on $p$ and $q$. For more details on the Hochschild-Serre spectral sequence, we refer to [7, Chapter V, Section 10].
Equipped with natural subquotient topology on $E_{r}$, the Hochschild-Serre spectral sequence suggests that one may prove Vogan’s conjecture by establishing the Hausdorffness of $E_{r}$-terms for every $r\geq 0$. Along this strategy, we expect that the total cohomology space ${\mathrm{H}}^{\bullet}(\mathfrak{u},V)$ is Hausdorff for certain smooth representations $V$ whose underlying $(\mathfrak{g},K)$-module are not necessarily of finite length. The Hausdorffness of $E_{0}$-terms is clear from the construction, and our results show that the $E_{1}$-terms are indeed Hausdorff. However, for $r\geq 2$, the Hausdorffness of $E_{r}$-terms is still unresolved. We hope to investigate this in future research.
@@
Acknowledgment: The authors would like to thank Wei Xiao for suggesting the references [5, 6, 8] on the cubic Dirac operator. F. J. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via the grant SFB-TRR 358/1 2023-491392403. B. Sun was supported in part by National Key R & D Program of China No. 2022YFA1005300 and New Cornerstone Science Foundation.
@@ 4. Proof of the main theorem
We keep the notations and assumptions from previous sections. In particular, $G$ is a (possibly disconnected) compact Lie group with complexified Lie algebra $\mathfrak{g}$, $(\pi,V)$ is a smooth representation of $G$ and $\mathfrak{q}$ is a parabolic subalgebra of $\mathfrak{g}$ with nilpotent radical $\mathfrak{u}$. Use the notations in Section 3, $S=\wedge^{\bullet}\mathfrak{u}$, $\mathfrak{h}$ is a common Cartan subalgebra of $\mathfrak{l}$ and $\mathfrak{g}$, $D$ is the cubic Dirac operator with respect to $(\mathfrak{g},\mathfrak{l})$ and $\partial\in{\mathrm{End}}(V\otimes S)$ (resp. $\operatorname{d}\in{\mathrm{End}}(V\otimes S)$) are the boundary (resp. coboundary) map of Lie algebra homology (resp. cohomology). From now on, we view $D$ as an operator on $V\otimes S$.
@@
Let $\mathfrak{b}_{L}$ be the Borel subalgebra of $\mathfrak{l}$ associated with $\Delta_{\mathfrak{l}}^{+}$ and $T_{L}:=N_{L}(\mathfrak{b}_{L})$ be the Cartan subgroup of $L$. Recall that $T_{L}$ acts on $\mathfrak{h}^{*}$ by the coadjoint action, which preserves $\Delta_{\mathfrak{l}}^{+}$, and inner form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) is invariant under the coadjoint action. By the Cartan-Weyl theory (see [7, Chapter IV, Section 2]), there is a finite set $\{t_{1},\cdots,t_{k}\}$ of elements in the Cartan subgroup $T_{L}$ of $L$, such that as $\mathfrak{l}$-modules,